# arXiv:2602.13749v3[cond-mat.str-el] 02 Mar 2026

### v3

@@ Appendix D Thermodynamics analyzed by BCFT
where $Q$ is total charge of fermions (relative to the ground state), $j$ and $j_{f}$ are integer or half-integers, and $n_{Q},n_{s},n_{f}$ denote the charge, spin, flavor index in conformal tower, respectively, which are non-negative integers.
For example, when $n=2$, one can obtain several lowest energy levels (in unit $El/\pi$) as following,
@@
The leading irrelevant operator in this case, $\mathbf{J}_{-1}\phi_{j}$ contribute to the impurity free energy in the following scaling rule,
@@
Specially, for $n=2$, the possible leading operator in anti-Hermitian sector comes one-level higher descendant, e.g., $\sim i\mathbf{J}_{-2}\cdot\phi$, with the scaling dimension $\Delta=5/2$. This non-Hermitian irrelevant operator leads to corrections of thermodynamic quantities at higher temperatures. In terms of the impurity specific heat, it brings about a correction,
@@
where $a,b$ are non-universal dimensionless parameters. Notably, from the NRG calculation in Fig.6(d)(e), we can see that the non-Hermicity has little effect on the specific heat at low temperatures, which makes sense because we know from above that the non-Hermitian effect is a higher order correction. However, in the relatively high temperature regime, the numerically extracted specific heat indeed shows increasingly evident deviations from the Hermitian case, this reflects the effect of the non-Hermicity induced $bT^{3}$ term in Eq.(138). Thus, the NHNRG calculations are in consistence with the BCFT predictions.

### v4

@@ Appendix E Thermodynamics analyzed by BCFT
where $Q$ is the total charge of fermions (relative to the ground state), $j$ and $j_{f}$ are integers or half-integers, and $n_{Q},n_{s},n_{f}$ denote the charge, spin, flavor index in conformal tower, respectively, which are non-negative integers.
For example, when $n=2$, one can obtain several lowest energy levels (in unit $El/\pi$) as follows,
@@
The leading irrelevant operator in this case, $\mathbf{J}_{-1}\phi_{j}$ contributes to the impurity free energy in the following scaling rule,
@@
Specially, for $n=2$, the possible leading operator in anti-Hermitian sector comes from one-level higher descendant, e.g., $\sim i\mathbf{J}_{-2}\cdot\phi$, with the scaling dimension $\Delta=5/2$. This non-Hermitian irrelevant operator leads to corrections of thermodynamic quantities at higher temperatures. In terms of the impurity specific heat, it brings about a correction,
@@
where $a,b$ are non-universal dimensionless parameters. Notably, from the NRG calculation in Fig.6(d)(e), we can see that the non-Hermicity has little effect on the specific heat at low temperatures, which makes sense because we know from above that the non-Hermitian effect is a higher order correction. However, in the relatively high temperature regime, the numerically extracted specific heat indeed shows increasingly evident deviations from the Hermitian case, this reflects the effect of the non-Hermicity induced $bT^{3}$ term in Eq.(140). Thus, the NHNRG calculations are consistent with the BCFT predictions.